# SpikeGS: Reconstruct 3D scene captured by a fast-moving bio-inspired camera

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) has been proven to exhibit exceptional performance in reconstructing 3D scenes. However, the effectiveness of 3DGS heavily relies on sharp images, and fulfilling this requirement presents challenges in real-world scenarios particularly when utilizing fast-moving cameras. This limitation severely constrains the practical application of 3DGS and may compromise the feasibility of real-time reconstruction. To mitigate these challenges, we proposed Spike Gaussian Splatting (SpikeGS), the first framework that integrates the Bayer-pattern spike streams into the 3DGS pipeline to reconstruct 3D scenes captured by a fast-moving high temporal resolution color spike camera in one second. With accumulation rasterization, interval supervision, and a special designed pipeline, SpikeGS realizes continuous spatiotemporal perception while extracts detailed structure and texture from Bayer-pattern spike stream which is unstable and lacks details. Extensive experiments on both synthetic and real-world datasets demonstrate the superiority of SpikeGS compared with existing spike-based and deblurring 3D scene reconstruction methods. Codes and data will be released at anonymous link: https://spikegs.github.io.
@@ Introduction
Although many deblurring methods (Chen and Liu 2024; Zhao, Wang, and Liu 2024) attempt to address these challenges and free the constraint on camera motion, the limitations of traditional RGB camera sensors greatly hinder their ability to mitigate motion blur effectively. Bio-inspired cameras, characterized by their high temporal resolution, are considered promising solutions to address these issues at their core. Bio-inspired cameras such as event and spike (Huang et al. 2023) cameras have been widely applied in 3D reconstruction (Rudnev et al. 2023; Xiong et al. 2024; Zhu et al. 2024) of high-speed scenes with significant success. However, event-based camera approaches may lack adequate visual texture and frequently rely on supplementary traditional cameras for reconstruction (Qi et al. 2023), thereby constraining their applicability. Spike cameras, which capture rich visual textures, are more suitable for independent 3D reconstruction. However, prior solutions (Guo et al. 2024b; Zhu et al. 2024) have encountered difficulties in training and rendering efficiency, which posing an insurmountable challenge for real-time rendering and fast reconstruction. These methods also have not undergone extensive testing in real-world high-speed scenarios and have encountered challenges in reconstructing Bayer-pattern spike streams as the first-generation spike camera can only reproduce grayscale signals. Besides, as depicted in Figure LABEL:fig1, the Bayer-pattern spike stream lacks both geometric consistency and time-domain stability. The rapid reconstruction of precise structures and accurate irradiance from unstable spike streams poses a significant challenge.
@@
Bio-inspired sensors have shown their advantages in most computer version problems including novel views synthesizing. EventNeRF (Rudnev et al. 2023) and Ev-nerf (Hwang, Kim, and Kim 2023) synthesize the novel view in scenarios such as high-speed movements that would not be conceivable with a traditional camera by event supervision. Nonetheless, these works assume that event streams are temporally dense and low-noise which is inaccessible in practice. Robust e-NeRF (Low and Lee 2023) incorporates a more realistic event generation model to directly and robustly reconstruct NeRFs under various real-world conditions. DE-NeRF (Ma et al. 2023) and E2NeRF (Qi et al. 2023) extend event nerf to dynamic scenes and severely blurred images as NeRF researchers did. Recently, many researchers (Xiong et al. 2024; Wang et al. ) introduced 3D Gaussian Splatting framework for reconstructing appearance and geometry solely from event data, significantly improved the speed and quality of reconstruction. However, due to the absence of texture details in event data, all these approaches yield limited results. Spike cameras (Dong, Huang, and Tian 2021) can capture more texture information. NeRF based on spike cameras (Guo et al. 2024b; Zhu et al. 2024) achieves higher 3D scene reconstruction quality than event-based methods. However, the above methods still cannot quickly and efficiently reconstruct real-world 3D scenes while the camera moves at high speeds. They also have problems in dealing with Bayer-pattern spike streams from color spike camera.
@@
where ${A}(\mathbf{x},t)$ is the accumulation at time $t$, ${A_{\mathbf{x}}}(t)$ is the accumulation without reset before time $t$, ${L_{C}}(\mathbf{x},\tau)$ is the input current of a certain color at time $\tau$ (proportional to light intensity) and ${C\in\{R,G,B\}}$ denote the color determined by the Bayer-pattern color layout. Note that in a real camera, noise can interfere with the input current ${L_{C}}$ and threshold $\phi$. As shown in Fig. 3, each spike is read out at discrete time $nT,n\in\mathbb{N}$ ($T$ is a micro-second level). The output of the color spike camera is a spatial-temporal binary stream $S$ with $H\times W\times N$ size. The $H$ and $W$ are the height and width of the sensor, and $N$ is the temporal window size of the spike stream. Please refer to appendix for more details.
@@
Where $\lambda_{accu}$ and $\lambda_{in}$ are hyperparameters. $\lambda_{accu}$ starts at 0 in the early stages of training and gradually increases with the number of epochs. $\lambda_{in}$ decreases gradually with the number of epochs.
@@
Our code is based on 3DGS (Kerbl et al. 2023) and we train the models for $30000$ iterations on one NVIDIA A800 GPU with the same optimizer and hyper-parameters as 3DGS. The color spike camera shown in Fig. 2 is Spike M1K40-H4-Gen3 with $1000\times 1000$ resolution and the motor shown in Fig. 2 is DMR4315. We also provided the motor control program in the supplementary materials.
@@
We first compare our method with 3D-GS (Kerbl et al. 2023) and deblur-GS (Chen and Liu 2024) directly trained on the blurry images. Then, we compare our method against several 3D scene reconstruction methods for spike cameras, such as SpikeNeRF (Zhu et al. 2024) and Spike-NeRF (Guo et al. 2024b). Finally, given the absence of existing spike-based 3DGS methods, we compared our method with three spike-based image reconstruction methods with 3DGS: TFI (Zhu et al. 2019)+3DGS, TFP (Zhu et al. 2019)+3DGS and Spk2imgNet (Zhao et al. 2021) +3DGS. We conduct experiments using proposed SpikeGS datasets to evaluate our method’s performance on both synthetic and real-world scenes. For synthetic datasets, we evaluate the synthesized novel view in terms of Peak Singal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). For real-world datasets, we evaluate the synthesized novel view using a set of No-Reference Image Quality Assessment (NR-IQA) metrics since we don’t have ground truth. The adopted metrics include NIQE (Mittal, Soundararajan, and Bovik 2012) and IL-NIQE (Zhang, Zhang, and Bovik 2015).
@@
Accumulation Loss and Interval Loss. We also conduct novel view synthesis experiments to investigate the effect of interval and accumulation on the same dataset. Tab. 3 shows that both interval and accumulation loss significantly enhance novel view synthesis quality. Moreover, the impact of accumulation loss remains consistent across both low-speed and high-speed scenarios while the influence of interval loss is notably more pronounced in high-speed scenarios than in low-speed scenarios. These results align with our theory: Interval recovers geometric details using minimal numbers of spikes so it’s more sensitive to speed.

### v4

@@ Abstract
3D Gaussian Splatting (3DGS) has been proven to exhibit exceptional performance in reconstructing 3D scenes. However, the effectiveness of 3DGS heavily relies on sharp images, and fulfilling this requirement presents challenges in real-world scenarios particularly when utilizing fast-moving cameras. This limitation severely constrains the practical application of 3DGS and may compromise the feasibility of real-time reconstruction. To mitigate these challenges, we proposed Spike Gaussian Splatting (SpikeGS), the first framework that integrates the Bayer-pattern spike streams into the 3DGS pipeline to reconstruct 3D scenes captured by a fast-moving high temporal resolution color spike camera in one second. With accumulation rasterization, interval supervision, and a special designed pipeline, SpikeGS realizes continuous spatiotemporal perception while extracts detailed structure and texture from Bayer-pattern spike stream which is unstable and lacks details. Extensive experiments on both synthetic and real-world datasets demonstrate the superiority of SpikeGS compared with existing spike-based and deblurring 3D scene reconstruction methods.
Codehttps://spikegs.github.io.
@@ Introduction
Figure 1: Brief introduction of SpikeGS. Given high temporal resolution but texture lacking spike stream captured by a fast-moving spike camera in one second, our method can reconstruct a sharp radiance field in minutes and achieves clear real-time rendering results.
Although many deblurring methods (Chen and Liu 2024; Zhao, Wang, and Liu 2024) attempt to address these challenges and free the constraint on camera motion, the limitations of traditional RGB camera sensors greatly hinder their ability to mitigate motion blur effectively. Bio-inspired cameras, characterized by their high temporal resolution, are considered promising solutions to address these issues at their core. Bio-inspired cameras such as event and spike (Huang et al. 2023) cameras have been widely applied in 3D reconstruction (Rudnev et al. 2023; Xiong et al. 2024; Zhu et al. 2024) of high-speed scenes with significant success. However, event-based camera approaches may lack adequate visual texture and frequently rely on supplementary traditional cameras for reconstruction (Qi et al. 2023), thereby constraining their applicability. Spike cameras, which capture rich visual textures, are more suitable for independent 3D reconstruction. However, prior solutions (Guo et al. 2024b; Zhu et al. 2024) have encountered difficulties in training and rendering efficiency, which posing an insurmountable challenge for real-time rendering and fast reconstruction. These methods also have not undergone extensive testing in real-world high-speed scenarios and have encountered challenges in reconstructing Bayer-pattern spike streams as the first-generation spike camera can only reproduce grayscale signals. Besides, as depicted in Figure 1, the Bayer-pattern spike stream lacks both geometric consistency and time-domain stability. The rapid reconstruction of precise structures and accurate irradiance from unstable spike streams poses a significant challenge.
@@
Bio-inspired sensors have shown their advantages in most computer version problems including novel views synthesizing. EventNeRF (Rudnev et al. 2023) and Ev-nerf (Hwang, Kim, and Kim 2023) synthesize the novel view in scenarios such as high-speed movements that would not be conceivable with a traditional camera by event supervision. Nonetheless, these works assume that event streams are temporally dense and low-noise which is inaccessible in practice. Robust e-NeRF (Low and Lee 2023) incorporates a more realistic event generation model to directly and robustly reconstruct NeRFs under various real-world conditions. DE-NeRF (Ma et al. 2023) and E2NeRF (Qi et al. 2023) extend event nerf to dynamic scenes and severely blurred images as NeRF researchers did. Recently, many researchers (Xiong et al. 2024) introduced 3D Gaussian Splatting framework for reconstructing appearance and geometry solely from event data, significantly improved the speed and quality of reconstruction. However, due to the absence of texture details in event data, all these approaches yield limited results. Spike cameras (Dong, Huang, and Tian 2021) can capture more texture information. NeRF based on spike cameras (Guo et al. 2024b; Zhu et al. 2024) achieves higher 3D scene reconstruction quality than event-based methods. However, the above methods still cannot quickly and efficiently reconstruct real-world 3D scenes while the camera moves at high speeds. They also have problems in dealing with Bayer-pattern spike streams from color spike camera.
@@
where ${A}(\mathbf{x},t)$ is the accumulation at time $t$, ${A_{\mathbf{x}}}(t)$ is the accumulation without reset before time $t$, ${L_{C}}(\mathbf{x},\tau)$ is the input current of a certain color at time $\tau$ (proportional to light intensity) and ${C\in\{R,G,B\}}$ denote the color determined by the Bayer-pattern color layout. Note that in a real camera, noise can interfere with the input current ${L_{C}}$ and threshold $\phi$. As shown in Fig. 4, each spike is read out at discrete time $nT,n\in\mathbb{N}$ ($T$ is a micro-second level). The output of the color spike camera is a spatial-temporal binary stream $S$ with $H\times W\times N$ size.Please refer to appendix for more details.
@@
Where $\lambda_{accu}$ and $\lambda_{in}$ are hyperparameters. $\lambda_{accu}$ starts at 0 in the early stages of training and gradually increases with the number of epochs.
@@
Our code is based on 3DGS (Kerbl et al. 2023) and we train the models for $30000$ iterations on one NVIDIA A800 GPU with the same optimizer and hyper-parameters as 3DGS. The color spike camera shown in Fig. 3 is Spike M1K40-H4-Gen3 with $1000\times 1000$ resolution and the motor shown in Fig. 3 is DMR4315.
@@
We first compare our method with 3D-GS (Kerbl et al. 2023) and deblur-GS (Chen and Liu 2024) directly trained on the blurry images. Then, we compare our method against several 3D scene reconstruction methods for spike cameras, such as SpikeNeRF (Zhu et al. 2024) and Spike-NeRF (Guo et al. 2024b). Finally, given the absence of existing spike-based 3DGS methods, we compared our method with three spike-based image reconstruction methods with 3DGS: TFI (Zhu et al. 2019)+3DGS, TFP (Zhu et al. 2019)+3DGS and Spk2imgNet (Zhao et al. 2021) +3DGS. We conduct experiments using proposed SpikeGS datasets to evaluate our method’s performance on both synthetic and real-world scenes. For synthetic datasets, we evaluate the synthesized novel view in terms of PSNR, SSIM, and LPIPS. For real-world datasets, we evaluate the synthesized novel view using a set of No-Reference Image Quality Assessment (NR-IQA) metrics since we don’t have ground truth. The adopted metrics include NIQE (Mittal, Soundararajan, and Bovik 2012) and IL-NIQE (Zhang, Zhang, and Bovik 2015).
@@
Accumulation Loss and Interval Loss. We also conduct novel view synthesis experiments to investigate the effect of interval and accumulation on the same dataset. Tab. 3 shows that both interval and accumulation loss significantly enhance novel view synthesis quality. Moreover, the impact of accumulation loss remains consistent across both low-speed and high-speed scenarios while the influence of interval loss is notably more pronounced in high-speed scenarios.